OpenStudy (anonymous):

What is the slope of the line x = -5? Undefined 0 -5 5

OpenStudy (anonymous):

the line is vertical a vertical line does not have a slope, go with "undefined"

OpenStudy (anonymous):

Thank you! Would you mind helping me with the rest of my exam?

OpenStudy (solomonzelman):

Yes, it is undefined. You can choose any vertical line, and plug in any 2 points into the slope formula, and you will surely find that it is UNDEFINED.

OpenStudy (anonymous):

lol don't mention "exam" people will get all upset

OpenStudy (anonymous):

lol okay then. But can you help?

OpenStudy (solomonzelman):

with ?

OpenStudy (anonymous):

What is the slope of the line between (3, -4) and (-2, 1)? 1 2 -2 -1

OpenStudy (solomonzelman):

Considering points (a,b) and (c,d) , \(\LARGE\color{blue}{ \rm{slope}= \frac{b-d}{a-c} }\)

OpenStudy (anonymous):

\[\frac{1-(-4)}{-2-3}=\frac{-4-1}{3-(-2)}=-1\]

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

in your case, you have (3,-4) and (-2,1) So, \(\LARGE\color{blue}{ \rm{slope}= \frac{(-4)-(1)}{(3)-(-2)}=\frac{-5}{5}=-1 }\)

OpenStudy (solomonzelman):

Satelline is fast.

OpenStudy (anonymous):

Thank you both! I have 4 more, if you don't mind?

OpenStudy (anonymous):

Alex is saving money because she would like to purchase the new video gaming system that everyone has been talking about. After two weeks of saving, she had $47. After 24 weeks of saving, she had $430. Approximately how many dollars per week did Alex save on average between weeks 2 and 24? $15.96 per week $17.41 per week $17.91 per week $19.88 per week

OpenStudy (solomonzelman):

24 weeks after the 2 weeks, or 24 weeks and 2 weeks are within them ?

OpenStudy (solomonzelman):

saying is it 6 or 24 altogether.

OpenStudy (solomonzelman):

I mean 26 nor 24

OpenStudy (anonymous):

I think it's excluding the 1st 2 weeks and it just means how much she saved between weeks 2-24

OpenStudy (solomonzelman):

*or

OpenStudy (solomonzelman):

weeks 1-2 she saves 47 AND weeks 1-24 (all weeks) she saves 430

OpenStudy (solomonzelman):

the average between 2-24 weeks, I would think that the weeks it is asking about is 3,4,5,...23,24. 430 - 47 = 383 (is how much she saves starting after the week 2) 383 ÷ 22 = average starting from the 3rd week, and till the end of 24th

OpenStudy (solomonzelman):

383 ÷ 22 = 17.4090909090909090.... ≈ 17.41

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

I couldn't copy this one in so I had to attach it as a jpg